# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ E.2 Detailed derivation of the mean firing rate in the quenched-noise approximation
Appendix F Alternative figures for different values of the model parameters
@@
In Figs. 16,17 we show our numerical and analytical estimations of ${v_{\rm R}}^{1/2}$ versus ${\mu_{\rm I}}$ (for two values of ${\tau_{\rm c}}$) and versus ${\tau_{\rm c}}$ (for two values of ${\mu_{\rm I}}$, with constant ${v_{\rm c}}$), respectively. We there report as well our NIS-corrected estimations for the rate variance. We have not used such estimations, since they present a relatively larger statistical error (than their ${r_{\rm R}}$ counterpart). We observe that, as happens for ${r_{\rm R}}$, and rather remarkably, for low and high enough values of ${\tau_{\rm c}}$, the finite-NIS-corrected estimations $\tilde{v_{\rm R}}^{1/2}$ become, respectively, compatible with the Poisson-Moreno and the quenched-noise approximations.
@@
Figure 15: Mean readout firing rate ${r_{\rm R}}$ vs ${\tau_{\rm c}}$: comparison between theoretical values and numerical estimations. Panels (A,B,C,D) correspond, respectively, to the different values of ${\mu_{\rm I}}=32.65,35.0,44.45,50.35$. Green points and error-bars are fixed-NIS numerical estimations; blue diamonds and error-bars are the finite-NIS-corrected estimations; the dotted black line is the Moreno extrapolated function in [17], while the pink dashed line is the quenched-noise approximation, Eq. (9).
Figure 16: Standard deviation of the readout firing rate ${v_{\rm R}}^{1/2}$ vs ${\mu_{\rm I}}$: comparison between theoretical values and numerical estimations. Panels A,B correspond, respectively, to the different values of ${\tau_{\rm c}}=0.1,0.72$. Green points and error-bars are fixed-NIS numerical estimations; blue diamonds and error-bars are the finite-NIS-corrected estimations; the dotted black line is the Poisson-Moreno approximation ${v_{\rm R}}={r_{\rm R}}_{\rm M}/{T_{\rm c}}$, while the gray dashed line is the Poisson-quenched approximation ${v_{\rm R}}={r_{\rm R}}_{\text{q}}/{T_{\rm c}}$; the pink dashed line fork correspond to the upper and lower-bounds of the quenched-noise approximation, Eq. (13). The rest of the simulation parameters are: ${\sigma_{\rm c}}=30.593$, ${\alpha_{\rm V}}=0.9$, ${\tau_{\rm m}}=0.005$, $w=0.45$, ${v_{\rm V}}=9.8$.
Figure 17: Standard deviation of the readout firing rate ${v_{\rm R}}^{1/2}$ vs ${\tau_{\rm c}}$: comparison between theoretical values and numerical estimations. Panels A,B correspond, respectively, to two different values of ${\mu_{\rm I}}=41.5,50.35$. The rest of the simulation parameters are: ${\sigma_{\rm I}}=4.760$, ${\sigma_{\rm c}}=0.361$, ${\alpha_{\rm V}}=0.9$, ${\tau_{\rm m}}=0.005$, $w=0.45$, $vvV=9.8$.
Figure 18: Readout SNR in the unit time interval, as in Fig. 8 but for different values of the tuning curves ${\mu_{\rm I}}^{(\pm)}$. The simulation parameters are $(\mu_{I}^{(+)},\mu_{I}^{(-)})=(32.650,41.500)$, $w=0.45$, ${v_{\rm V}}=9.8$.
Figure 19: Temporal distribution function $g(t)$, as in Fig. 6 but for a different value of ${\mu_{\rm I}}=56.25$.
Appendix G Analytical expression for the rate variance in the quenched-noise approximation
G.1 Sketch of the derivation of the temporal distribution function in the quenched-noise approximation

### v2

@@ E.2 Derivation of the mean firing rate in the quenched-noise approximation
E.3 Relation with asymptotic integral expansions
Taylor-expanding $r$ around $y_{0}$ within the integrand in (46) would lead to the assymptotic integral expansion:
$I$ $=\left(\frac{1}{2\pi v^{2}}\right)^{1/2}\int_{\Delta}{\rm d}y\,\left(\sum_{m=0}^{\infty}(y-y_{0})^{m}\frac{r^{{}^{\prime}m}(y_{0})}{m!}\right){\cal N}(y;y_{0},v)=\sum_{m=1}^{\infty}\frac{r^{{}^{\prime}m}(y_{0})}{m!}\left<(y-y_{0})^{m}\right>_{{\cal N}(y;y_{0},v)}=$ (59)
$\simeq\sum_{n=0}^{\infty}\frac{r^{{}^{\prime}2n}(y_{0})}{(2n)!}v^{n}$ (60)
where $r^{{}^{\prime}p}$ means the $p$-th order derivative. The last equation is an approximation, because we have approximated the integrating interval $[0,\infty]$ by [,][$\infty$,$\infty$][ - ∞ , ∞ ]. We notice that the standard Laplace approximation for the integral is the 00-th order of the assymptotic expansion in $v$. Our expression, Eq. (52) coincides with the assymptotic expansion up to first order in $v$. Indeed, it is easy to expand Eq. 52 to the first order in $v$, one finds:
$I$ $=r(y_{0})+\frac{1}{2}vr^{\prime\prime}(y_{0})+O[v^{3}]$ (61)
which trivially coincides with (60) up to $O[v]$. This confirms again that our expression, eq. 52 is a refinement of the standard Laplace expression.
E.4 Formal mathematical proof
Let $g$ and $h$ be two functions, twice continuous and differentiable in an interval $\Delta$ in the positive real line, that can also be unbounded. Now, suppose that $g$ has a single absolute maximum at $x_{0}\in\Delta$, and it is $g^{\prime\prime}_{0}:=g^{\prime\prime}(x_{0})<0$. We want to evaluate the integral
$I_{M}=\int_{\Delta}{\rm d}x\,e^{Mg(x)+h(x)}$ (62)
The standard saddle point approximation is:
$I^{\rm(std)}_{M}=e^{Mg_{0}+h_{0}}\left(\frac{2\pi}{M|g_{0}^{\prime\prime}|}\right)^{1/2}$ (63)
and one can easily proof that
$\lim_{M\to\infty}\frac{I_{M}}{I^{\rm(std)}_{M}}=1.$ (64)
Let us consider the alternative approximation:
$I^{\rm(alt)}_{M}=e^{Mg_{0}+h_{0}}\left(\frac{2\pi}{M|g_{0}^{\prime\prime}|-h_{0}^{\prime\prime}}\right)^{1/2}\exp{\left(\frac{1}{2}\frac{M^{-1}{h^{\prime}_{0}}^{2}}{|g_{0}^{\prime\prime}|-M^{-1}h_{0}^{\prime\prime}}\right)}.$ (65)
It is easy to proof that it also satisfies:
$\lim_{M\to\infty}\frac{I_{M}}{I^{\rm(alt)}_{M}}=1$ (66)
but it is less immediate to see that it is actually a better approximation than $I^{\rm(std)}_{M}$. Indeed, it is:
$\lim_{M\to\infty}\frac{I_{M}-I^{\rm(alt)}_{M}}{I_{M}-I^{\rm(std)}_{M}}=1$ (67)
and not $<1$, as one could initially have expected! The reason is that the alternative approximation is a correction for large, but not infinite values of $M$. Rather, we want to proof that the quotient above is $\leq 1$ for large enough $M$. The proposition to be demonstrated is the following. Given $g,h,\Delta,x_{0}$, there is some $\tilde{M}$ above which it is
$\left|\frac{I_{M}-I^{\rm(alt)}_{M}}{I_{M}-I^{\rm(std)}_{M}}\right|\leq 1\qquad\forall M>\tilde{M}.$ (68)
Let us see how this could be proven. We need some preliminary propositions.
Lower and upper bounds. We follow the demonstration in [81]. First, we are able to write analytical expressions for lower and upper bounds to $I_{M}$, that will be referred to as $I_{M}^{{\rm lb}(\epsilon)},I_{M}^{{\rm ub}(\epsilon)}$ respectively, both depending on a real parameter $\epsilon>0$. For the lower bound, one increases by $\epsilon$ the curvatures of $g$ and $f$, in such a way that $\forall\epsilon>0,\exists\,\delta>0$ such that $Mg(x)+h(x)\geq Mg_{0}-M(|g^{\prime\prime}(x)|+\epsilon)\delta_{x}^{2}/2+h_{0}+h^{\prime}_{0}\delta_{x}+(h^{\prime\prime}_{0}-\epsilon)\delta_{x}^{2}/2$ for all $x$ in the interval $|\delta_{x}|<\delta$. As a consequence:
$I_{M}$ $\geq\int_{|\delta_{x}|<\delta}{\rm d}x\,e^{Mg(x)+h(x)}\geq I_{M}^{{\rm lb}(\epsilon)}$ (69)
$I_{M}^{{\rm lb}(\epsilon)}$ $:=e^{Mg_{0}+h_{0}}\int_{-\delta M^{1/2}}^{\delta M^{1/2}}{\rm d}y\,e^{-y^{2}(|g_{0}^{\prime\prime}|+\epsilon)/2+y^{2}M^{-1}(h^{\prime\prime}_{0}-\epsilon)/2+yM^{-1/2}h_{0}^{\prime}}.$ (70)
For the upper bound, we decrease the curvature by a quantity $\epsilon>0$, which is small enough so that the increased curvature is still lower than zero, $-g^{\prime\prime}_{0}+\epsilon<0$. By continuity of $g^{\prime\prime}$ and the Taylor theorem, for each of such $\epsilon$’s $\exists\,\delta>0$ such that
$Mg(x)+h(x)\leq M(g_{0}-(|g^{\prime\prime}_{0}|-\epsilon)\delta_{x}^{2}/2)+h_{0}+h^{\prime}_{0}\delta_{x}+(h^{\prime\prime}_{0}+\epsilon)\delta_{x}^{2}/2$ (71)
for all $x$ such that $|\delta_{x}|<\delta$. Now, let $M^{*}$ be a value for which the integral
$\tilde{I}:=\int_{\Delta}{\rm d}x\,e^{M^{*}g(x)+h(x)}<\infty$ (72)
is finite. Hence, given $\epsilon$ and $\delta$, there exist a $\eta>0$ such that $Mg(x)\leq M(g_{0}-\eta)$ in all the points $|\delta_{x}|>\delta$ outside the interval $(-\delta+x_{0},x_{0}+\delta)$. Therefore,
$I_{M}$ $=$ $\int_{|\delta_{x}|>\delta}{\rm d}x\,e^{Mg(x)+h(x)}+$ $\int_{|\delta_{x}|<\delta}{\rm d}x\,e^{Mg(x)+h(x)}\leq$ (73)
$\leq$ $\int_{|\delta_{x}|>\delta}{\rm d}x\,e^{M^{*}g(x)}e^{(M-M^{*})(g_{0}-\eta)+h(x)}+$ $\int_{|\delta_{x}|<\delta}{\rm d}x\,e^{Mg(x)+h(x)}\leq$ (74)
$\leq$ $e^{(M-M^{*})(g_{0}-\eta)}\tilde{I}+$ $\int_{|\delta_{x}|<\delta}{\rm d}x\,e^{Mg(x)+h(x)}\leq$ (75)
$\leq$ $e^{(M-M^{*})(g_{0}-\eta)}\tilde{I}+$ $e^{Mg_{0}+h_{0}}\int_{|\delta_{x}|<\delta}{\rm d}x\,e^{-M(|g^{\prime\prime}_{0}|-\epsilon)\delta_{x}^{2}/2+\delta_{x}h^{\prime}_{0}+(h^{\prime\prime}_{0}+\epsilon)\delta_{x}^{2}/2}\leq$ (76)
$\leq$ $e^{(M-M^{*})(g_{0}-\eta)}\tilde{I}+$ $e^{Mg_{0}+h_{0}}\int_{-\infty}^{\infty}{\rm d}x\,e^{-M(|g^{\prime\prime}_{0}|-\epsilon)\delta_{x}^{2}/2+h_{0}+\delta_{x}h^{\prime}_{0}+(h^{\prime\prime}_{0}+\epsilon)\delta_{x}^{2}/2},$ (77)
or, using Gaussian formulae:
$I_{M}\leq I_{M}^{{\rm ub}(\epsilon)}=e^{(M-M^{*})(g_{0}-\eta)}\tilde{I}+e^{Mg_{0}+h_{0}}\left(\frac{2\pi}{M(|g^{\prime\prime}_{0}|-\epsilon)-(h_{0}^{\prime\prime}+\epsilon)}\right)^{1/2}\exp{\left(\frac{1}{2}\frac{{h^{\prime}_{0}}^{2}}{M(|g_{0}^{\prime\prime}|-\epsilon)-(h^{\prime\prime}_{0}+\epsilon)}\right)}.$ (78)
The strategy to demonstrate the proposition in (68) is to examine, for large $M$, the following lower and upper bounds
$q_{M}^{{\rm u}(\epsilon)}$ $:=\frac{I_{M}^{{\rm ub}(\epsilon)}-I^{\rm(alt)}_{M}}{I_{M}^{{\rm lb}(\epsilon)}-I^{\rm(std)}_{M}}$ (79)
$q_{M}^{{\rm l}(\epsilon)}$ $:=\frac{I_{M}^{{\rm lb}(\epsilon)}-I^{\rm(alt)}_{M}}{I_{M}^{{\rm ub}(\epsilon)}-I^{\rm(std)}_{M}}$ (80)
for the ratio:
$q_{M}^{{\rm l}(\epsilon)}\leq\frac{I_{M}-I^{\rm(alt)}_{M}}{I_{M}-I^{\rm(std)}_{M}}\leq q_{M}^{{\rm u}(\epsilon)}\qquad\forall\epsilon.$ (81)
We demonstrate that $q_{M}^{{\rm u}(\epsilon)}\leq 1$ and that $q_{M}^{{\rm l}(\epsilon)}\geq-1$ for large enough $M$. Let us consider $q_{M}^{{\rm u}(\epsilon)}$. It is easy to see that:
$q_{M}^{{\rm u}(\epsilon)}=\frac{e^{-\eta M}\,u\,\tilde{I}+\left(\frac{2\pi}{M(|g^{\prime\prime}_{0}|-\epsilon)-(h_{0}^{\prime\prime}+\epsilon)}\right)^{1/2}e^{\frac{1}{2}\frac{{h^{\prime}_{0}}^{2}}{M(|g_{0}^{\prime\prime}|-\epsilon)-(h_{0}^{\prime\prime}+\epsilon)}}-\left(\frac{2\pi}{M|g_{0}^{\prime\prime}|-h_{0}^{\prime\prime}}\right)^{1/2}e^{\frac{1}{2}\frac{{h^{\prime}_{0}}^{2}}{M|g_{0}^{\prime\prime}|-h_{0}^{\prime\prime}}}}{\int_{-\delta M^{1/2}}^{\delta M^{1/2}}{\rm d}y\,e^{-\frac{y^{2}}{2}(|g_{0}^{\prime\prime}|+\epsilon-M^{-1}(h_{0}^{\prime\prime}-\epsilon))+yM^{-1/2}h_{0}^{\prime}}-\left(\frac{2\pi}{M|g_{0}^{\prime\prime}|}\right)^{1/2}.}$ (82)
where $u:=\exp[-M^{*}(g_{0}-\eta)-h_{0}]$. It is instructive to see what happens if we take first the $M\to\infty$ limit of this expression:
$\lim_{M\to\infty}\frac{I_{M}^{{\rm ub}(\epsilon)}-I^{\rm(alt)}_{M}}{I_{M}^{{\rm lb}(\epsilon)}-I^{\rm(std)}_{M}}=\frac{(|g_{0}^{\prime\prime}|-\epsilon)^{-1/2}-(|g_{0}^{\prime\prime}|)^{-1/2}}{(|g_{0}^{\prime\prime}|+\epsilon)^{-1/2}-(|g_{0}^{\prime\prime}|)^{-1/2}}.$ (83)
Our best estimators for $I_{M}$ are recovered in the $\lim_{\epsilon\to 0}$, which gives
$\lim_{\epsilon\to 0}\lim_{M\to\infty}\frac{I_{M}^{{\rm ub}(\epsilon)}-I^{\rm(alt)}_{M}}{I_{M}^{{\rm lb}(\epsilon)}-I^{\rm(std)}_{M}}=1.$ (84)
Now, it is clear what is happening: in the $\lim_{M\to\infty}$ our corrections to the standard Laplace method (i.e., the differences between $I^{({\rm alt})}$ and $I^{({\rm std})}$) vanish. Let us now consider the ratio in (82), and suppose that $M$ is so large that one can neglect the exponentially decaying term $\exp{(-\eta M)}$ in the numerator, and that we can substitute the integral in the denominator by the same integral with the ,$\infty$,$\infty$- ∞ , ∞ limits (the differences between both decay very fast as the complementary error function $\rm erfc$ in $\delta M^{1/2}$, whose leading term is $\sim\delta M^{1/2}\exp(\delta^{2}M)$). Indeed, these are the only two terms decaying exponentially in $M$. We have:
$\frac{I_{M}^{{\rm ub}(\epsilon)}-I^{\rm(alt)}_{M}}{I_{M}^{{\rm lb}(\epsilon)}-I^{\rm(std)}_{M}}\simeq\frac{\left(\frac{2\pi}{M(|g^{\prime\prime}_{0}|-\epsilon)-(h_{0}^{\prime\prime}+\epsilon)}\right)^{1/2}e^{\frac{1}{2}\frac{{h^{\prime}_{0}}^{2}}{M(|g_{0}^{\prime\prime}|-\epsilon)-(h_{0}^{\prime\prime}+\epsilon)}}-\left(\frac{2\pi}{M|g_{0}^{\prime\prime}|-h_{0}^{\prime\prime}}\right)^{1/2}e^{\frac{1}{2}\frac{{h^{\prime}_{0}}^{2}}{M|g_{0}^{\prime\prime}|-h_{0}^{\prime\prime}}}}{\left(\frac{2\pi}{M(|g^{\prime\prime}_{0}|+\epsilon)-(h_{0}^{\prime\prime}-\epsilon)}\right)^{1/2}e^{\frac{1}{2}\frac{{h^{\prime}_{0}}^{2}}{M(|g_{0}^{\prime\prime}|+\epsilon)-(h_{0}^{\prime\prime}-\epsilon)}}-\left(\frac{2\pi}{M|g_{0}^{\prime\prime}|}\right)^{1/2}}\qquad\text{large $M$}.$ (85)
Mind that, as far as $h^{\prime}_{0}\neq 0$ or $h^{\prime\prime}_{0}\neq 0$, this ratio is lower than one in absolute value, because the difference between both terms in the numerator is lower in absolute value than the difference between both terms in the denominator (and it would vanish for $\epsilon=0$). Notice that, given $M$, $\epsilon$ cannot be arbitrary low for (85) to be a good approximation of (82): it has to be $\epsilon$ (and, consequently, $\delta$ and $\eta$ in (82)) large enough so that we can neglect $\exp{(-\eta M)}$ and substitute $\int_{-\delta M^{1/2}}^{\delta M^{1/2}}$ by $\int_{-\infty}^{\infty}$. The larger $M$, the lower we can take $\epsilon$, and the better is the estimation in (85). However, it is not necessary that $\epsilon$ is very small. The only thing we need is $M$ large enough for the upper bound (85) to be $\leq 1$. The same argument can be applied to $q_{M}^{{\rm l}(\epsilon)}\geq-1$, and this leads us to Eq. (68).
E.5 Validity of the analytic expression for the mean readout rate in the quenched-noise approximation
We have seen that the condition in which this equation is derived is low enough ${v_{\rm c}}$. This condition can be actually seen to be satisfied in the simulations corresponding to the main analysis of this article: the 1st simulation for constant ${v_{\rm c}}$ presented in Secs. III,IV. For our derivation to hold (see above in this appendix), we need to neglect powers of $O[(\delta_{\mu})^{3}]$ of $\ln{r_{\rm w}}$ in an interval of the order of few multiples of ${v_{\rm c}}^{1/2}$ around ${\mu_{\rm I}}$. Indeed, this is what happens even in the less favorable case (that of the higher ${v_{\rm c}}({\alpha_{\rm V}}=0.9)\simeq 25.4994$ for ${\alpha_{\rm V}}=0.9$) in the 1st simulation, see Fig. 22. As a consequence, the analytic expression for ${r_{\rm q}}$ in Eq. (54) coincides with the numeric estimation of the integral in the general expression for ${r_{\rm q}}$, Eq. (45), as we illustrate in Fig. (5-B).
We remind that in the 3rd simulation (analysed in Appendix F), the RO neuron properties are analysed as a function of ${\sigma_{\rm V}}$, for constant ${\tau_{\rm c}}$, which implies varying ${v_{\rm c}}$. For the highest values of ${\sigma_{\rm V}}$, the value of ${v_{\rm c}}$ lies outside the validity limits of the derivation of Eq. (54), and this equation begins to differ from the numerical integration of Eq. (45) (see Fig. 23). This example illustrates the fact that the validity limit of the analytical expression for the mean rate in the quenched-noise approximation, ${r_{\rm q}}$ in Eq. (54) has a validity limit (low ${v_{\rm c}}$) that, in general, differs from that of the quenched approximation itself (Eq. (45) for the mean and Eq. (87) for the temporal distribution function), which is ${\tau_{\rm c}}\gg{\tau_{\rm m}}$. For the highest values of ${\sigma_{\rm V}}$, hence of ${v_{\rm c}}$ in Fig. 23, the analytic expression for ${r_{\rm q}}$ is no longer valid, while the quenched noise approximation still holds (since it is ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.005$).
Figure 22: Necessary conditions for the demonstration of the quenched-noise approximation for the rate mean in Eq. (10). The derivation holds in the extent to which terms of $O[(\mu-{\mu_{\rm I}})^{3}]$ of $\ln{r_{\rm w}}$ in Eq. (50) are negligible in intervals of few standard deviations ${v_{\rm c}}^{1/2}$ of ${\mu_{\rm I}}$ around its mean ${\mu_{\rm I}}$ (outside which the contribution of the integral (46) is negligible). We plot the different terms in the exponent $f(\mu)$ in Eq. (50), with ${\mu_{\rm I}}=45$ and the parameters of the article simulations for ${\alpha_{\rm V}}=0.9$. Indeed, within an interval of length $6{v_{\rm c}}^{1/2}$ around ${\mu_{\rm I}}$, the $\ln{r_{\rm w}}$ is well approximated by its first two orders in $\delta_{\mu}$. The big circle and its error-bar indicate the mean ${\mu_{\rm I}}$ and (twice the) standard deviation ${v_{\rm c}}^{1/2}$ of the normal distribution.
Appendix F SNR behavior varying ${\sigma_{\rm c}}$ at fixed ${\tau_{\rm c}}$
As a complementary analysis of that of Sec. IV, we have analysed the behavior of the SNR in the NERM versus the value of ${\sigma_{\rm V}}$ (consequently ${\sigma_{\rm c}}$) for a fixed value of ${\tau_{\rm c}}$. The simulation parameters of the corresponding simulation (the 3rd simulation, see Appendix C) are specified in Table 6. The results are summarised in Fig. 23.
The salient features of the figure are the following. First, the mean readout rate enhancement $\Delta{r_{\rm R}}$ in panel A is, again statistically consistent, at least for low enough ${\sigma_{\rm V}}$, with the quenched-noise approximation. The discrepancies for large ${\sigma_{\rm V}}$ are attributed to the finite-NIS effect, as we observed in Sec. IV. The analytical expression for the quenched-noise effect, Eq. (10) coincides with its numerical counterpart, Eq. (9), for low enough ${\sigma_{\rm V}}$ (i.e., low enough ${v_{\rm c}}$), as expected from Sec. E.2. The same arguments apply to the stimulus-averaged rate standard deviation in panel B. As a consequence of the underestimation of the variance due to the finite-NIS effect, the SNR is overestimated for large ${\sigma_{\rm V}}$, and departs from the analytical prediction. Overall, this case illustrates the non-monotonic dependence of the SNR on ${\sigma_{\rm c}}$, for which the analytical solution provides an interpretable explanation (see Sec. IV).
In this case, and differently from the so called 1st simulation, the finite-NIS reduction algorithm does not help us to construct a better estimation of the SNR. The reason is that the underestimation of the rate variance ${v_{\rm R}}$ due do the finite-NIS effect is more severe for this simulation parameters. As we wrote above, unfortunately the finite-NIS reduction algorithm does not provide us with an accurate estimation for the rate variance, given the used values of ${n_{\rm is}}$ and ${T_{\rm c}}$, as it does for the rate mean. Therefore, although the finite-NIS estimation for the rate mean is closer to the theoretical ${r_{\rm q}}$, the finite-NIS correction does not help much to construct a better numerical estimation of the SNR. We therefore conclude that it is the theoretical curve in Fig. 23-C to be correct, and the points corresponding to the numerical simulations to be an overestimation of the SNR due to the finite-NIS effect.
Figure 23: Mean rate enhancement, rate standard deviation and SNR, as a function of ${\sigma_{\rm V}}$. As in Fig. 9 but for a fixed value of ${\tau_{\rm c}}=0.1$, and varying ${\sigma_{\rm V}}$. See the rest of model parameters in Table 6.
Appendix G Additional figures for alternative values of the model parameters
@@
In Figs. 25,26 we show our numerical and analytical estimations of ${v_{\rm R}}^{1/2}$ versus ${\mu_{\rm I}}$ (for two values of ${\tau_{\rm c}}$) and versus ${\tau_{\rm c}}$ (for two values of ${\mu_{\rm I}}$, with constant ${v_{\rm c}}$), respectively. We there report as well our NIS-corrected estimations for the rate variance. We have not used such estimations, since they present a relatively larger statistical error (than their ${r_{\rm R}}$ counterpart). We observe that, as happens for ${r_{\rm R}}$, and rather remarkably, for low and high enough values of ${\tau_{\rm c}}$, the finite-NIS-corrected estimations $\tilde{v_{\rm R}}^{1/2}$ become, respectively, compatible with the Poisson-Moreno and the quenched-noise approximations.
@@
Figure 24: Mean readout firing rate ${r_{\rm R}}$ vs ${\tau_{\rm c}}$: comparison between theoretical values and numerical estimations. Panels (A,B,C,D) correspond, respectively, to the different values of ${\mu_{\rm I}}=32.65,35.0,44.45,50.35$. Green points and error-bars are fixed-NIS numerical estimations; blue diamonds and error-bars are the finite-NIS-corrected estimations; the dotted black line is the Moreno extrapolated function in [17], while the pink dashed line is the quenched-noise approximation, Eq. (9). See the rest of model parameters in Table 4.
Figure 25: Standard deviation of the readout firing rate ${v_{\rm R}}^{1/2}$ vs ${\mu_{\rm I}}$: comparison between theoretical values and numerical estimations. Panels A,B correspond, respectively, to the different values of ${\tau_{\rm c}}=0.1,0.72$. Green points and error-bars are fixed-NIS numerical estimations; blue diamonds and error-bars are the finite-NIS-corrected estimations; the dotted black line is the Poisson-Moreno approximation ${v_{\rm R}}={r_{\rm R}}_{\rm M}/{T_{\rm c}}$, while the gray dashed line is the Poisson-quenched approximation ${v_{\rm R}}={r_{\rm R}}_{\text{q}}/{T_{\rm c}}$; the pink dashed line fork correspond to the upper and lower-bounds of the quenched-noise approximation, Eq. (13). See the rest of model parameters in Table 4.
Figure 26: Standard deviation of the readout firing rate ${v_{\rm R}}^{1/2}$ vs ${\tau_{\rm c}}$: comparison between theoretical values and numerical estimations. Panels A,B correspond, respectively, to two different values of ${\mu_{\rm I}}=41.5,50.35$. See the rest of model parameters in Table 4.
Figure 27: Readout SNR in the unit time interval, as in Fig. 9 but for different values of the tuning curves ${\mu_{\rm I}}^{(\pm)}$. The tuning curves are $(\mu_{I}^{(+)},\mu_{I}^{(-)})=(35.6,50.35)$. The thin dotted lines are the encoding signal-to-noise ratio in the unit time interval, Eq. (8). See the rest of model parameters in Table 4.
Figure 28: Temporal distribution function $g(t)$, as in Fig. 7 but for a different value of ${\mu_{\rm I}}=56.25$. See the rest of model parameters in Table 4.
Appendix H Analytical expression for the rate variance in the quenched-noise approximation
H.1 Sketch of the derivation of the temporal distribution function in the quenched-noise approximation